OpenStudy (lina17):

solve: -2(2x+5)-3=-3(x-1)

OpenStudy (lina17):

@rsadhvika

OpenStudy (lina17):

help please

OpenStudy (rsadhvika):

Notice that the equation has lot of "negative" signs

OpenStudy (rsadhvika):

I don't like "negative" signs

OpenStudy (rsadhvika):

what should we do to get rid of negative signs ?

OpenStudy (lina17):

multiply

OpenStudy (lina17):

?

OpenStudy (rsadhvika):

Yes, multiply "-1" through out and the equation becomes : \[ 2(2x+5)+3=3(x-1)\]

OpenStudy (rsadhvika):

Any idea on next step ?

OpenStudy (lina17):

no clue

OpenStudy (rsadhvika):

Maybe, distribute

OpenStudy (rsadhvika):

2(2x+5) = ?

OpenStudy (rsadhvika):

3(x-1) = ?

OpenStudy (lina17):

it would be

OpenStudy (lina17):

(4x+10)-3=(3x-3

OpenStudy (lina17):

@rsadhvika

OpenStudy (rsadhvika):

Yes, \[4x+10-3 = 3x-3\]

OpenStudy (lina17):

k

OpenStudy (rsadhvika):

group the like terms 10 and -3

OpenStudy (rsadhvika):

10 - 3 = ?

OpenStudy (lina17):

4x-7=3x-3

OpenStudy (rsadhvika):

\[4x+7 = 3x-3\]

OpenStudy (rsadhvika):

it is +7 ok

OpenStudy (lina17):

k

OpenStudy (rsadhvika):

what next

OpenStudy (rsadhvika):

our goal here is to solve x

OpenStudy (lina17):

we add 4x+7

OpenStudy (rsadhvika):

what good is that ?

OpenStudy (rsadhvika):

\[4x+7 = 3x-3\]

OpenStudy (rsadhvika):

I'd subtract 3x both sides

OpenStudy (lina17):

oh ok

OpenStudy (rsadhvika):

that gets rid of "x" terms on the right hand side

OpenStudy (rsadhvika):

\[x+7= -3\]

OpenStudy (lina17):

x+7=-3

OpenStudy (rsadhvika):

Next subtract 7 both sides and you're done!

OpenStudy (rsadhvika):

\[x = -3-7\]

OpenStudy (rsadhvika):

\[x = -10\]

OpenStudy (lina17):

oh ok thank you so much

OpenStudy (rsadhvika):

yw